OpenStudy (anonymous):

I need help on how to to square out the cube in this problem. (x+2)^3= -27

myininaya (myininaya):

ok you want to take the cube root to get rid of the cube (x+2)=(-27)^(1/3)

OpenStudy (anonymous):

Oh, then what would I do from there?

myininaya (myininaya):

x+2=[(-3)(-3)(-3)]^(1/3) x+2=-3 x=-3-2=-5 check: x=-5 (x+2)^3=-27 (-5+2)^3=-27 (-3)^3=-27 -27=-27

OpenStudy (anonymous):

Oh =O I see now. Thank you ^^